Consider the following aggregate expenditure model of the Canadian economy operating with given wages and other factor prices, price level, interest rates, exchange rates, and expectations:
C = 50 + 0.8YD I = 400 G = 500 T = 0.3Y X = 650 IM = 0.36Y
where C is consumption (the 0.8 term represents the marginal propensity to consume) YD is disposable income, I is investment, G is government spending on goods and services, T is the total value of taxes net of transfers (the 0.3 term represents the net tax rate on national income), X is exports, and IM is imports (the 0.36 term represents the marginal propensity to import).

Now suppose that the government decides to use its spending power to restore national income to its original level. By how much must the government increase G to restore the original level of national income? What will happen to the government’s budget balance?
1
Expert's answer
2015-02-27T12:13:33-0500
C = 50 + 0.8YD
I = 400
G = 500
T = 0.3Y
X = 650
IM = 0.36Y
(b) Y1 = $1520.9
(c) Y2 = $1420.9
(d) If the government decides to use its spending power to restore national income to its original level, and as G = 500, the government must increase G by 100 to restore the original level of national income.
The new government’s budget balance will be:
BS = 0.3*1520.9 - 600 = -$143.7, so it will decrease by 70, so the budget deficit will increase.